OpenStudy (timinthor):

Laura fills a bucket with water at a rate of 6 L/min. At the same time, Sarah empties a bucket holding 30 L of water at a rate of 9 L/min. Let l represent the number of liters of water and let t represent time in minutes. The system models this situation. How many minutes will it take for the buckets to hold equal amounts of water, and how much water will that be? l = 6t l = 30 – 9t It will take ___ minutes for both buckets to hold equal amounts of water. They will each hold ___ liters.

Vocaloid (vocaloid):

hint: set the equations equal to each other and solve for t 6t = 30 - 9t know where to go from here?

OpenStudy (timinthor):

I think so, im going to try

OpenStudy (timinthor):

6t=30-9t -6t -6t 30=3t

OpenStudy (timinthor):

am i right so far?

Vocaloid (vocaloid):

hm, not quite. what is -9t - 6t?

OpenStudy (timinthor):

Ohhhh i forgot the -9 not a positive 9 my bad, it would be -15t right?

Vocaloid (vocaloid):

yeah, so 30 = 15t, which means that t = 2 minutes

Vocaloid (vocaloid):

to find the volume, we just use one of the original equations L = 6t and plug in t = 2

OpenStudy (timinthor):

L=6(2) so it would be 12

Vocaloid (vocaloid):

yup! that's our final answer. 2 minutes and 12 liters

OpenStudy (timinthor):

Okay! Thank you! ill give a medal

OpenStudy (timinthor):

Yay! I got 100% on the rest of the test too! thx!